# PPR: Enhancing Dodging Attacks while Maintaining Impersonation Attacks on Face Recognition Systems

### v2

@@ Abstract
Adversarial Attacks on Face Recognition (FR) encompass two types: impersonation attacks and evasion attacks. We observe that achieving a successful impersonation attack on FR does not necessarily ensure a successful dodging attack on FR in the black-box setting. Introducing a novel attack method named Pre-training Pruning Restoration Attack (PPR), we aim to enhance the performance of dodging attacks whilst avoiding the degradation of impersonation attacks. Our method employs adversarial example pruning, enabling a portion of adversarial perturbations to be set to zero, while tending to maintain the attack performance. By utilizing adversarial example pruning, we can prune the pre-trained adversarial examples and selectively free up certain adversarial perturbations. Thereafter, we embed adversarial perturbations in the pruned area, which enhances the dodging performance of the adversarial face examples. The effectiveness of our proposed attack method is demonstrated through our experimental results, showcasing its superior performance.
@@ 1 Introduction
Figure 1: An example of comparison between our proposed attack (blue) and the traditional impersonation attack (red) in the black-box setting. The letters ‘D’ and ‘S’ signify that the FR model classifies the two face images as belonging to different identities and belonging to the same identity, respectively.
Thanks to the ceaseless advancements in deep learning, FR has made remarkable strides in achieving exceptional levels of performance(Schroff et al., 2015)(Wang et al., 2018)(Deng et al., 2022). However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security.
Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models. Consequently, several research endeavors have been directed towards this realm.
@@
A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness(Qiu et al., 2020)(Yang et al., 2021)(Cherepanova et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023), transferability(Zhong & Deng, 2021)(Zhou et al., 2023b)(Zhou et al., 2023a), and physical attack capability(Yin et al., 2021)(Yang et al., 2023)(Li et al., 2023b). These efforts contribute to enhancing the effectiveness of existing attacks using face adversarial examples. However, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using the impersonation attacks.
Unlike image classification, FR is an open-set task. Predicting the class probability of identities in real-world deployment is an extremely challenging task. In practical FR applications, we extract embeddings from two face images using the FR model. Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity. If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities. Based the measurement of FR, there exists the samples that can be classified as two different identities in theory. We denote these samples as multi-identity samples. Our experiments have verified the existence of multi-identity samples on common-used face dataset (refer to Appendix A).
In impersonation attacks, the objective is to deceive the FR model into recognizing adversarial face examples as a predefined identity, which requires that the distance between the embeddings of the adversarial face example and the victim image be lower than the threshold. On the other hand, in dodging attacks, the aim is to trick the FR model into recognizing the adversarial face example as identities other than the attacker, which means that the distance between the embeddings of the adversarial face example and the attacker image should be greater than the threshold. The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR in theory. Our experiment confirms that the majority of adversarial face examples crafted by traditional impersonation attacks, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting (see Appendix B).
In practical application scenarios, criminals tend to generate adversarial examples with their own faces to deceive FR models into recognizing them as victims for impersonation attacks, while avoiding being recognized as criminals themselves for dodging attacks to escape legal prosecution. This necessitates the generation of adversarial examples that can perform both impersonation and dodging attacks.
To improve the dodging performance of the adversarial face examples without degrading the impersonation performance, we propose a novel attack method, termed as Pre-training Pruning Restoration Attack (PPR). Specifically, inspired by the pruning methods in model compression, we prune the pre-trained adversarial face examples to sparsify the adversarial face examples. Subsequently, we add adversarial perturbations that favor dodging attacks on the pruned regions to improve the dodging performance of the adversarial face examples (i.e. the operation of Restoration). After pruning and restoration, the impersonation performance of the adversarial face examples can be maintained and the dodging performance of the adversarial face examples can be improved.
Figure 1 illustrates a comparison between our proposed attack and the traditional impersonation attack. The adversarial face example generated by our proposed attack method successfully evades identification as the attacker in contrast with the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker. The adversarial face example generated by our proposed attack method successfully evades identification as the attacker, while the adversarial face example crafted by the traditional impersonation attack method is identified as the attacker. Note that both the adversarial face examples crafted by our proposed attack method and the traditional impersonation attack method are capable of being identified as the victim image.
@@
  • We verify the universality of multi-identity samples among adversarial face examples crafted by traditional impersonation attacks and demonstrate that a successful impersonation attack does not necessarily equate to a successful dodging attack on FR systems.
  • We propose a novel adversarial attack method called PPR based on adversarial example pruning. By pruning the adversarial face examples and subsequently introducing adversarial perturbations favoring dodging attacks, we effectively enhance the dodging performance. The approach allows us to maintain the impersonation performance while significantly improving the dodging performance of the adversarial face examples.
  • Extensive experiments demonstrate that our proposed method achieves superior performance compared to the baseline adversarial attack methods.
2 Related Works
@@ 2.1 Adversarial Attacks
The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images for the purpose of deceiving machine learning systems and causing them to make mistakes(Szegedy et al., 2014)(Goodfellow et al., 2015). The existence of adversarial examples poses a significant threat to the security of current machine learning systems. Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems(Long et al., 2022)(Zhang et al., 2022)(Liang et al., 2023)(Wei et al., 2023)(Lu et al., 2023).
To improve the performance of black-box adversarial attacks, DI is proposed by (Xie et al., 2019) to randomly transform adversarial examples in each iteration. VMI-FGSM, proposed by (Wang & He, 2021), employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance. SSA, as proposed by (Long et al., 2022), transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them. (Wang et al., 2023b) propose SIA. SIA applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculations. (Wang et al., 2023a) propose BSR. BSR divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation.
2.2 Face Recognition
FR is an open-set task, meaning we cannot predict the class of a user upon deploying an FR model. As a result, relying solely on the maximum probability value of the output of the neural network to determine the class the of a user becomes impractical. Instead, when deploying the FR model, we extract the embeddings of two images and compute the similarity between them. If the similarity exceeds a predefined threshold, we conclude that the two images belong to the same identity; otherwise, they belong to different identities.
Recent research on FR primarily focuses on reducing the intra-class distance and increasing the inter-class distance of embeddings extracted by FR models. (Schroff et al., 2015) introduce FaceNet that employs triplet loss to map face images into a space where the distance represents the dissimilarity between them. (Wang et al., 2018) develop CosFace, which utilizes the large-margin cosine loss to enhance FR model performance. Additionally, (Deng et al., 2022) introduce ArcFace that incorporates the additive angular margin loss to train FR models that extract more discriminative features.
@@ 2.3 Adversarial Attacks on Face Recognition
Adversarial attacks on FR can be mainly classified into two categories: restricted attacks and unrestricted attacks.
Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g. $L_{p}$ bound). Because FR belongs to the open-set task, the embedding-level loss is more suitable. Therefore, (Zhong & Deng, 2019) proposed FIM that use embedding-level loss to craft the adversarial face examples. (Zhong & Deng, 2021) proposed DFANet that performs dropout on the feature maps of the convolutional layers to improve the transferability. (Yang et al., 2021) propose TIP-IM to improve the face encryption. TIP-IM use MMD(Borgwardt et al., 2006) loss to improve the quality of encrypted face images and use the greedy insertion to select the optimal victim images. (Zhou et al., 2023b) propose BPFA to improve the transferability of adversarial attacks on FR by adding beneficial perturbations(Wen & Itti, ) on the features maps of the FR models to obtain effect that is similar to hard samples augmentation.
The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined bound. The unrestricted adversarial attacks on FR mainly focus on physical attacks(Xiao et al., 2021)(Yang et al., 2023)(Li et al., 2023a), attribute editing(Jia et al., 2022) and generating adversarial examples based on makeup transfer(Yin et al., 2021)(Hu et al., 2022)(Shamshad et al., 2023).
@@ 3.1 Problem Formulation
In the context of our study, $\mathcal{F}^{vct}(x)$ denotes the FR model used by the victim to extract the embedding from a face image $x$. We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively. The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ into misclassifying $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$. To be more specific, the objective can be expressed as follows:
where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation. By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$. In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
Due to the differing objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types. When it comes to impersonation attacks on FR, the ASR can be computed as:
where the value of $t^{d}$ represents the threshold for dodging attacks.
@@ 3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it impossible to optimize objectives as stated in Equation 1 and Equation 2 directly. To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversaries that can be transferred to the victim model for an effective attack(Yuan et al., 2022)(Zhang et al., 2022)(Naseer et al., 2023)(Wang et al., 2023b)(Wang et al., 2023a).
@@
where $\phi(x)$ represents the operation that normalizes $x$. $x^{adv}$ is the adversarial example which is initialized with the same value as $x^{s}$. Dodging attacks can be formulated with the following loss function:
@@
As the FR task is an open-set task, it is not possible to predict the classes of users during the practical deployment of the FR model. Therefore, we need to compare the similarity between two face images in order to discern whether they depict the same identity or not. Based on the identification method in face recognition, multi-identity samples exist theoretically. Our experiments verify the existence of such samples among benign face images. The existence of multi-identity samples raises a question:
To this end, we generated face adversarial samples using the traditional impersonation attack and evaluate its dodging ASR (refer to Appendix B). Our experiments demonstrate that the majority of adversarial face examples crafted by traditional impersonation attack cannot achieve a successful dodging attack in the black-box setting.
Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this could lead to legal consequences. Hence, it is crucial to research attack techniques that can execute both impersonation and dodging attacks simultaneously. Traditional impersonation attacks on FR systems have shown a remarkably high level of impersonation ASR in black-box settings. Therefore, our objective is to enhance the dodging performance while maintaining the impersonation effectiveness of traditional impersonation attacks.
To accomplish this objective, a straightforward approach is to generate adversarial face examples by employing a Lagrangian attack strategy that utilizes the following loss function:
@@
However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance leading to subpar performance (See Section 4.2). If we alliviete the trade-off, we can achieve a better dodging performance while maintaining the impersonation performance.
3.3 Pre-training Pruning Restoration Attack
To accomplish this objective, a straightforward approach is to fine-tune the adversarial face examples generated by the Lagrangian attack with a lower $\lambda$ value in order to enhance the performance of dodging attacks. However, this method does not enhance the dodging attack performance without compromising the impersonation attack performance (see Fine-tuning in Table 4). We contend that this issue arises because the newly introduced adversarial perturbation that favors dodging attacks ends up disrupting the pre-existing adversarial perturbation patterns. While it may improve the performance of dodging attacks, it inevitably diminishes the performance of impersonation attacks. To address this, we introduce new adversarial perturbations favoring dodging attacks in regions where the original perturbations have not been added. However, identifying suitable areas for these new perturbations is challenging due to their scarcity. Therefore, drawing inspiration from prior works in model pruning(Ma et al., 2023)(Frantar & Alistarh, 2023), we propose incorporating pruning into the realm of adversarial face examples. The aim is to free up some model parameters while minimizing the decrease in the performance of impersonation attacks, thereby creating space to introduce perturbations that facilitate dodging attacks.
In order to prune the adversarial perturbation, our initial step is to assess its importance by identifying and sparsifying the less important regions. To estimate this importance, we propose utilizing the magnitude of the adversarial perturbation as a measure. A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification.
Let the adversarial examples be $x^{adv}$. The formula to calculate the importance can be express as:

### v3

# Rethinking Impersonation and Dodging Attacks on Face Recognition Systems

@@ Abstract
Face Recognition (FR) systems can be easily deceived by adversarial examples that manipulate benign face images through imperceptible perturbations. Adversarial attacks on FR encompass two types: impersonation (targeted) attacks and dodging (untargeted) attacks. Previous methods often achieve a successful impersonation attack on FR; However, it does not necessarily guarantee a successful dodging attack on FR in the black-box setting. In this paper, our key insight is that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously. To this end, we propose a novel attack method termed as Adversarial Pruning (Adv-Pruning), to fine-tune existing adversarial examples to enhance their dodging capabilities while preserving their impersonation capabilities. Adv-Pruning consists of Priming, Pruning, and Restoration stages. Concretely, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations, identifying and releasing those with minimal impact on absolute model output variances. Then, Biased Gradient Adaptation is presented to adapt the adversarial examples to traverse the decision boundaries of both the attacker and victim by adding perturbations favoring dodging attacks on the vacated regions, preserving the prioritized features of the original perturbations while boosting dodging performance. As a result, we can maintain the impersonation capabilities of original adversarial examples while effectively enhancing dodging capabilities. Comprehensive experiments demonstrate the superiority of our method compared with state-of-the-art adversarial attacks.
@@ 1 Introduction
Figure 1: Top: previous methods that achieve a successful impersonation attack on FR cannot guarantee a successful dodging attack on FR in the black-box setting. In contrast, we present Adv-Pruning, including Priming, Pruning, and Restoration Stages, to perform both impersonation and dodging attacks simultaneously. Bottom (left): natural Multi-identity Samples (MS). Bottom (right): the dodging Attack Success Rate (%) between the previous methods and Adv-Pruning on multiple models.
@@
Thanks to the ceaseless advancements in deep learning, Face Recognition (FR) has achieved exceptional performance [41; 49; 9; 1; 2; 26]. However, the vulnerability of existing FR models to adversarial attacks poses a significant threat to their security. Hence, there is an urgent need to enhance the performance of adversarial face examples to expose more blind spots in FR models. As a result, several research endeavors have been directed towards this realm. A multitude of adversarial attacks have been developed to create adversarial face examples with characteristics such as stealthiness [40; 62; 7; 19; 42], transferability [70; 29; 72; 71], and physical attack capability [64; 63; 28]. These efforts contribute to enhancing the effectiveness of adversarial attacks on FR. Nevertheless, these studies primarily concentrate on bolstering either impersonation attacks or dodging attacks, overlooking the exploration of the effectiveness of dodging attacks when crafting adversarial face examples using impersonation attacks.
In real-world deployment contexts, individuals with malicious intent are prone to creating adversarial face examples incorporating their own facial features to manipulate FR systems to mistakenly identify them as pre-defined victims during impersonation attacks. Concurrently, the individuals strive to evade accurate identification as perpetrators, thereby circumventing detection and preventing legal accountability. This requires the creation of adversarial examples capable of executing both impersonation and dodging attacks simultaneously. In the realm of adversarial attacks on image classification, a successful impersonation attack typically implies a successful dodging attack. However, FR is an open-set task [49; 9], which is quite different from image classification. In the real-world deployment of FR systems, accurately predicting the class probability of identities presents an extreme challenge. Therefore, we extract embeddings from two face images using the FR model. Subsequently, the distance between the two embeddings is used to determine whether the images belong to the same identity. If the distance falls below a predefined threshold, the two images are recognized as belonging to the same identity; otherwise, they are classified as different identities. Based on the measurement of FR, there are two decision boundaries for each FR model when crafting adversarial examples as shown in Fig. 1. As a result, there exists the natural samples that can be classified as two different identities in theory. We denote these samples as multi-identity samples (refer to Section 4.4).
The existence of multi-identity samples implies that a successful impersonation attack on FR does not necessarily guarantee a successful dodging attack on FR. Existing research indicates that an adversarial sample was located near the decision boundary [18; 4]. Suppose we generate adversarial face examples using previous methods. In the white-box setting, both the structures and parameters of the victim models are known, enabling the generation of adversarial face examples that can cross the decision boundaries of both attacker and victim, as shown in Fig. 1. However, in the black-box setting, the decision boundaries of black-box models differ from those of the surrogate models. Consequently, adversarial examples generated on the surrogate model lie near the decision boundary of the victim, preventing them from crossing the decision boundary of the attacker. As such, the majority of adversarial face examples crafted by previous methods, which can successfully perform impersonation attacks, fail to perform dodging attacks in the black-box setting.
Figure 2: Overview of our Adv-Pruning attack framework, which consists of Priming, Pruning, and Restoration stages. (a) During the Priming stage, we optimize the adversarial examples to ensure they have sufficient attack performance. (b) In the Pruning stage, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations. Subsequently, we sparsify the adversarial perturbations based on the quantified priorities. (c) In the Restoration stage, we present Biased Gradient Adaptation to introduces gradient perturbations biased to dodging attacks on the sparsified regions.
In this paper, we propose a novel attack method, termed as Adversarial Pruning (Adv-Pruning). In the realm of adversarial attacks on FR, previous impersonation methods have achieved a significant level of sophistication. However, there remains a pressing need to bolster the efficacy of adversarial face examples in dodging attacks. Consequently, our research is directed towards enhancing the dodging attack performance of adversarial face examples while maintaining the impersonation attack performance. Specifically, we introduce an attack consisting of three stages: Priming, Pruning, and Restoration. In the Priming stage, we optimize the adversarial examples to ensure adequate attack potential. In the Pruning stage, with considering the pruning concept in model compression, we propose Adversarial Priority Quantification to measure the region-wise priority of original adversarial perturbations using an priority measure which is directly proportional to the supremum of the absolute model output variances. After processing by Adversarial Priority Quantification, we prune the adversarial face examples to free up less prioritized adversarial perturbations. In the Restoration stage, we propose Biased Gradient Adaptation to add biased gradient perturbations favoring dodging attacks on the pruned regions to adapt the adversarial face examples into the space that can be classified as the victim while remaining unidentifiable as the attacker, thereby enhancing the dodging performance of the adversarial face examples without compromising the prioritized features of original adversarial perturbations. As illustrated in the top of Fig. 1, after undergoing these stages, the adversarial face example generated by our proposed method can successfully traverse the decision boundaries of both the attacker and victim of the black-box model, achieving successful black-box impersonation and dodging attacks.
@@
  • We offer a new perspective for adversarial attacks on FR models that the generation of adversarial examples should perform both impersonation and dodging attacks simultaneously. To the best of our knowledge, this is the first work that studies the universality of multi-identity samples among adversarial face examples crafted by impersonation attacks.
  • We propose a novel adversarial attack method called Adversarial Pruning (Adv-Pruning). Adversarial Priority Quantification is presented to quantify the priority of the adversarial perturbations with minimal impact on absolute model output variances. Biased Gradient Adaptation is designed to adapt the adversarial examples to traverse both the decision boundaries of attacker and victim using biased gradients.
  • Extensive experiments demonstrate that our proposed method achieves superior performance compared to the state-of-the-art adversarial attack methods. Moreover, our presented method could be plugged into various FR systems and adversarial attack methods.
2 Related Work
@@ 2.1 Adversarial Attacks
The primary objective of adversarial attacks is to introduce imperceptible perturbations to benign images to deceive machine learning systems and cause them to make mistakes [47; 14]. The existence of adversarial examples poses a significant threat to the security of current machine learning systems. Lots of efforts have been dedicated to researching adversarial attacks in order to enhance the robustness of these systems [33; 69; 30; 55; 34; 43; 12; 73; 68; 37]. To improve the performance of black-box adversarial attacks, DI [60] applies random transformations to adversarial examples in each iteration to achieve a data augmentation effect. VMI-FGSM [53] employs gradient variance to stabilize the updating process of adversarial examples, boosting the black-box performance. SSA [33] transforms adversarial examples into the frequency domain and uses spectrum transformation to augment them. SIA [54] applies a random image transformation to each image block, generating a varied collection of images that are then employed for gradient calculation. BSR [51] divides the input image into multiple blocks, subsequently shuffling and rotating these blocks in a random manner, creating a collection of new images for the purpose of gradient calculation. DA [16] utilizes dispersion amplification to enhance the multi-task attack capability of adversarial attacks. Despite their gratifying progress, these studies neglect the consideration of pruning adversarial examples through introducing pruning methods into the realm of adversarial attacks. In our research, we propose a novel pruning method capable of identifying and freeing up the adversarial perturbations with minimal impact on absolute model output variances, thereby sparsifying regions for adding adversarial perturbations with the aim of dodging capabilities improvement.
@@ 2.2 Adversarial Attacks on Face Recognition
Based on the restriction of the adversarial perturbations, adversarial attacks on FR can be classified into two categories: restricted attacks [32; 61; 10; 74; 35; 31; 5] and unrestricted attacks [3; 6; 65; 8; 44; 46; 48; 56; 57]. Restricted attacks on FR are the attacks that generate adversarial examples in a restricted bound (e.g. $L_{p}$ bound). To enhance the transferability of adversarial attacks on FR, [70] propose DFANet, which applies dropout on the feature maps of the convolutional layers to achieve ensemble-like effects. In addition, [72] introduces BPFA, which further improves the transferability of adversarial attacks on FR by incorporating beneficial perturbations [58] on the feature maps of the FR models, resulting in hard model augmentation effects. [29] leverages extra information from FR-correlated tasks and uses a multi-task optimization framework to enhance the transferability of crafted adversarial examples. The unrestricted adversarial attacks on FR are the attacks that generate adversarial examples without the restriction of a predefined perturbation bound. They mainly focus on physical attacks [59; 63; 28], attribute editing [40; 23] and generating adversarial examples based on makeup transfer [64; 19; 42]. The existing literature on both restricted and unrestricted adversarial attacks on FR systems has successfully enhanced the performance of these attacks. Nevertheless, it remains under-explored in the correlation between impersonation and dodging attacks. This paper elegantly addresses this by investigating the correlation between impersonation and dodging attacks and introducing a novel attack method that bolsters the dodging capabilities while preserving the impersonation capabilities of previous methods.
@@ 3.1 Problem Formulation
Let $\mathcal{F}^{vct}(x)$ denote the FR model used by the victim to extract the embedding from a face image $x$. We refer to $x^{s}$ and $x^{t}$ as the attacker and victim images, respectively. The objective of the impersonation attacks explored in our research is to manipulate $\mathcal{F}^{vct}$ in order to misclassify $x^{adv}$ as $x^{t}$, while ensuring that $x^{adv}$ bears a close visual resemblance to $x^{s}$. By contrast, the objective of the dodging attacks proposed in this study is to render $\mathcal{F}^{vct}(x)$ unable to identify $x^{adv}$ as $x^{s}$, while simultaneously ensuring that $x^{adv}$ bears a visual resemblance to $x^{s}$. For the sake of clarity and conciseness, the detailed optimization objectives for both impersonation and dodging attacks are provided in the appendices.
Few works explore the correlation between impersonation and dodging attacks on FR. In the following, we delve into the correlation between these two types of attacks and propose a novel method to enhance dodging attacks while maintaining impersonation attacks. An overview of the proposed method is illustrated in Fig. 2. As depicted in Fig. 2, our proposed method is structured into three stages: Priming, Pruning, and Restoration. Through the sequential application of these stages, we are able to generate adversarial examples that exhibit a potent combination of impersonation and dodging attack capabilities.
@@ 3.2 Exploring the Impersonation and Dodging Attack on Face Recognition
In most cases, the victim model $\mathcal{F}^{vct}$ is not accessible to the attacker, making it extremely challenging to optimize the objectives for black-box attacks directly. To circumvent this issue, a common approach is to leverage a surrogate model $\mathcal{F}$ accessible to the attacker to generate adversarial examples that can be transferred to the victim model for an effective attack [66; 69; 39; 54; 51; 38; 11; 13; 27; 5].
@@
where $\phi(x)$ represents the operation that normalizes $x$. $x^{adv}$ is the adversarial example which is initialized with the same value as $x^{s}$. The loss function of dodging attacks can be formulated as:
@@
As the FR task is an open-set task, it is impractical to predict the classes of users during the practical deployment of the FR model. Therefore, we need to compare the distance between two face images to discern whether they depict the same identity or not. Based on the identification method in FR, multi-identity samples exist theoretically. Our experiments verify the existence of such samples among benign face images. The existence of multi-identity samples raises a question:
Does the success of an impersonation attack imply the success of dodging attacks on FR systems?
To this end, we generate adversarial face examples using the previous impersonation attack and evaluate its dodging Attack Success Rate (ASR). Our experiment confirms that the majority of adversarial examples crafted through previous methods, which are successful in performing impersonation attacks, fail to successfully execute dodging attacks in the black-box setting (see Section 4.4).
Nonetheless, in real-world adversarial attacks, attackers do not want the adversarial face examples to be recognized as themselves, as this may lead to legal consequences. Hence, it is crucial to research attack techniques that can execute both impersonation and dodging attacks simultaneously. Previous methods on FR systems have shown a remarkably high level of impersonation ASR in black-box settings. Therefore, our objective is to enhance the dodging performance while maintaining the impersonation effectiveness of previous attack methods.
To accomplish this objective, a straightforward approach is to generate adversarial face examples using a multi-task attack strategy. In the following, we will take the Lagrangian attack strategy as the example for its simplicity. The Lagrangian attack strategy utilizes the following loss function to craft adversarial examples:
@@
However, due to the conflict between the optimization between $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, there exists a trade-off between the performance of impersonation and dodging performance, leading to subpar performance (See Section 4.4). Suppose we can mitigate the trade-off, we will achieve a better dodging performance while maintaining the impersonation performance.
3.3 Adversarial Pruning Attack
To accomplish this objective, a straightforward approach is to fine-tune the adversarial face examples generated by the Lagrangian attack with a lower $\lambda$ value in order to enhance the performance of dodging attacks. However, this method does not enhance the dodging attack performance without compromising the impersonation attack performance (see Fine-tuning in Table 3). We contend that this issue arises because the newly introduced adversarial perturbation that favors dodging attacks ends up disrupting the prioritized features of existing adversarial perturbation. While it may improve the performance of dodging attacks, it inevitably diminishes the performance of impersonation attacks. To address this, we introduce new perturbations favoring dodging attacks in regions where original perturbations are not added. Nevertheless, identifying suitable areas for these new perturbations is challenging due to their scarcity. Therefore, we propose a novel pruning method to release less prioritized adversarial perturbations with minimal impact on the absolute model output variances, thereby creating space to introduce perturbations that facilitate dodging attacks.
Our proposed Adv-Pruning and be combined with various adversarial attacks. In the following, we will introduce our proposed Adv-Pruning based on Lagrangian attack in detail. In the Priming Stage, we utilize Eq. 3 as the Priming loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
$x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)\right)$ (4)
where $t$ is the iteration of the optimization process of adversarial examples, and $\beta$ is the step size when optimizing the adversarial face examples in the Priming stage, and $\prod\left(x\right)$ is the projection function that projects $x$ onto the $L_{p}$ norm bound.
Adversarial Priority Quantification. After completing the Priming Stage, we obtain an adversarial example with varying magnitudes of gradient perturbations across different regions. Following this, we proceed to the Pruning stage to process the crafted adversarial example. In order to prune the adversarial perturbation, our initial step is to assess its priority. To estimate this priority, we propose Adversarial Priority Quantification to quantify the priority of adversarial perturbations. Specifically, Adversarial Priority Quantification utilizes the magnitude of the adversarial perturbation as a measure. A lower magnitude implies a lesser impact on the performance of the adversarial examples generated after sparsification, as the supremum of absolute model output variances is directly proportional to the magnitude of the adversarial perturbations. The proof is in the appendices.
Let the adversarial examples be $x^{adv}$. The formula to calculate the priority can be expressed as: